OpenStudy (anonymous):

WILL MEDAL AND FAN: The perimeter of this rectangle can be calculated using the equation P = 2 × 18.5 + 2 × 7.2. Which other equation could be used to calculate the perimeter? A. P = (2 • 18.5) + 7.2 B. P = 2 (18.5 + 7.2) C. P = 2 × 18.5 + 7.2 D. P = 2 × 18.5 × 7.2

OpenStudy (anonymous):

@satellite73 @gorv @iGreen

OpenStudy (anonymous):

I forgot how to find another way to solve a rectangle

OpenStudy (anonymous):

-_- I am stupid

OpenStudy (igreen):

Option B. It multiplies 2 to both 18.5 and 7.2 just like in the original equation.

OpenStudy (igreen):

And they both equal 51.4.

OpenStudy (gorv):

take the 2 common

OpenStudy (gorv):

2(18.5+7.2)

OpenStudy (anonymous):

....

OpenStudy (anonymous):

I am confused

OpenStudy (igreen):

In 2(18.5 + 7.2), you will distribute the 2 like so: \((2 \times 18.5) + (2 \times 7.2)\)

OpenStudy (anonymous):

sorry :(

OpenStudy (igreen):

And that is the same as the original given equation P = 2 × 18.5 + 2 × 7.2.

OpenStudy (anonymous):

ok thanks